OpenStudy (anonymous):

Write a degree 3 polynomial with leading coefficient of 1 and zeros 5 3 -2

OpenStudy (campbell_st):

well you just need f(x) = (x -5)(x -3)(x + 2) then distribute and it will meet at the requirements

OpenStudy (anonymous):

would 3x^2-7x+9 be my answer then? I am just confused on how to distribute and need a refresher.

OpenStudy (campbell_st):

nope... you'll have \[(x^2 -8x + 15) \times ( x + 2) = x \times (x^2 -8x + 15) + 2\times (x^2 - 8x + 15)\] so simplify this